Pablo and faustino start runnin from the same location in the same direction. Pablo runs 8 miles per hour and faustino runs 6 miles per hour. After how long will they be 1/2 mile apart?
Answer by Alan3354(69443)

Pablo and faustino start runnin from the same location in the same direction. Pablo runs 8 miles per hour and faustino runs 6 miles per hour. After how long will they be 1/2 mile apart?
---------------
They separate at 8-6 = 2 mi/hr
---
(1/2)/2 = 1/4 hour